A potter’s wheel—a thick stone disk of radius 0.500 m and mass 100 kg—is freely rotating at 50.0 revs/min. The potter can stop the wheel in 6.00 s by pressing a wet rag against the rim and exerting a radially inward force of 70.0 N. Find the effective coefficient of kinetic friction between wheel and rag.



1
Expert's answer
2021-05-17T18:56:36-0400

"m = 100kg"

"R = 0.5m"

"N = 70"

"\omega_i = 50(\dfrac{rev}{min})(\dfrac{2\pi rad}{1 rev})(\dfrac{1min}{60s}) = 5.24rad/s"


"\omega_f = 0"

"t = 6s"

The moment of inertia of a solid disk,

"I = \dfrac{1}{2}mR^2"

After substituting the given values:

"I = \dfrac{1}{2}(100)(0.5)^2 = 12.5kgm^2"

Using the equation:

"\omega_f = \omega_i + \alpha t"


"\alpha = \dfrac{0-5.24}{6} = -0.873rad/s^2"


Now, using "\tau = I \alpha"

Putting values of "\alpha" and "I"

"\tau = (12.5)(-0.873) = -10.9"

The magnitude of torque due to kinetic frictional force is:

"\tau = fR"


"f = \dfrac{\tau}{R}"


"f = \dfrac{-10.9}{0.5} = -21.8"


The kinetic friction force "f = \mu_kN"

Hence,

"\mu_k = \dfrac{f}{N}"


"\mu_k= \dfrac{21.8}{70}= 0.311"